### v1

@@ 4.6 Evaluating the Performance of the Proposed Model in Comparison to Existing Research
Table 10: Result Analysis of existing study with the proposed system.
@@
MobileNetv2   R1 91.11%
VGG16   R11 81.48%
ResNet50   R11 74.07%
InceptionV3   R11 82.96%
Ensemble   R11 79.26%
VGG19-CBAM-Dense   R21 71.86%
DenseNet121-CBAM-Dense   R21 78.27%
EfficientNetB3-CBAM-Dense  R21 81.43%
MobileNetV2-CBAM-Dense  R21 74.07%
Xception-CBAM-Dense   R21 83.89%
@@ 5 Discussion
6 Conclusion
In conclusion, this study introduced a deep learning-based framework for effectively classifying skin lesions associated with monkeypox and non-monkeypox that achieved promising results. Through comprehensive experimentation and comparison with established pre-trained models, including VGG16, ResNet50, and DenseNet121, our research aimed to evaluate MpoxSLDNet’s effectiveness. The findings demonstrate MpoxSLDNet’s superior performance metrics, including accuracy, precision, recall, F1-score, and AUC, underscoring its potential as a reliable tool for early detection and categorization of monkeypox lesions. Table 8 demonstrates that the suggested MpoxSLDNet model showed superior performance compared to traditional pre-trained models, with higher accuracy, precision, recall, F1-score, and AUC metrics. The proposed MpoxSLDNet model achieved remarkable results, including an accuracy rate of 94.56%, a precision score of 0.94, a recall rate of 0.94, an F1-score of 0.94, and an AUC value of 0.94. In comparison, VGG16, DenseNet121, and ResNet50 obtained accuracy scores of 86.25%, 84.38%, and 67.19% respectively.Despite these promising outcomes, it is crucial to acknowledge inherent limitations in our study, particularly concerning the dataset’s representativeness of real-world clinical scenarios. Future research efforts should prioritize expanding the dataset to include a more diverse range of lesion variations and conducting rigorous clinical validation studies.
7 Limitation and Future Work

### v2

@@ IV-F Evaluating the Performance of the Proposed Model in Comparison to Existing Research
TABLE X: Result Analysis of existing study with the proposed system.
@@
MobileNetv2 [1] 91.11%
VGG16 [11] 81.48%
ResNet50 [11] 74.07%
InceptionV3 [11] 82.96%
Ensemble [11] 79.26%
VGG19-CBAM-Dense [20] 71.86%
DenseNet121-CBAM-Dense [20] 78.27%
EfficientNetB3-CBAM-Dense [20] 81.43%
MobileNetV2-CBAM-Dense [20] 74.07%
Xception-CBAM-Dense [20] 83.89%
@@ V Discussion
VI Conclusion
In conclusion, this study introduced a deep learning-based framework for effectively classifying skin lesions associated with monkeypox and non-monkeypox that achieved promising results. Through comprehensive experimentation and comparison with established pre-trained models, including VGG16, ResNet50, and DenseNet121, our research aimed to evaluate MpoxSLDNet’s effectiveness. The findings demonstrate MpoxSLDNet’s superior performance metrics, including accuracy, precision, recall, F1-score, and AUC, underscoring its potential as a reliable tool for early detection and categorization of monkeypox lesions. Table VIII demonstrates that the suggested MpoxSLDNet model showed superior performance compared to traditional pre-trained models, with higher accuracy, precision, recall, F1-score, and AUC metrics. The proposed MpoxSLDNet model achieved remarkable results, including an accuracy rate of 94.56%, a precision score of 0.94, a recall rate of 0.94, an F1-score of 0.94, and an AUC value of 0.94. In comparison, VGG16, DenseNet121, and ResNet50 obtained accuracy scores of 86.25%, 84.38%, and 67.19% respectively.Despite these promising outcomes, it is crucial to acknowledge inherent limitations in our study, particularly concerning the dataset’s representativeness of real-world clinical scenarios. Future research efforts should prioritize expanding the dataset to include a more diverse range of lesion variations and conducting rigorous clinical validation studies.
VII Limitation and Future Work